13. Allen Air Lines must liquidate some equipment that is being replaced. The equipment originally cost $22 million, of which 80% has been depreciated. The used equipment can be sold today for $7.7 million, and its tax rate is 40%. What is the equipment's after-tax net salvage value? Write out your answer completely. For example, 2 million should be entered as 2,000,000.

14. Although the Chen Company's milling machine is old, it is still in relatively good working order and would last for another 10 years. It is inefficient compared to modern standards, though, and so the company is considering replacing it. The new milling machine, at a cost of $40,000 delivered and installed, would also last for 10 years and would produce after-tax cash flows (labor savings and depreciation tax savings) of $9,500 per year. It would have zero salvage value at the end of its life. The Project cost of capital is 10%, and its marginal tax rate is 35%. Should Chen buy the new machine?

15. The Campbell Company is considering adding a robotic paint sprayer to its production line. The sprayer's base price is $1,080,000, and it would cost another $19,500 to install it. The machine falls into the MACRS 3-year class (the applicable MACRS depreciation rates are 33.33%, 44.45%, 14.81%, and 7.41%), and it would be sold after 3 years for $587,000. The machine would require an increase in net working capital (inventory) of $16,000. The sprayer would not change revenues, but it is expected to save the firm $446,000 per year in before-tax operating costs, mainly labor. Campbell's marginal tax rate is 30%. a. What is the Year 0 net cash flow? $ b. What are the net operating cash flows in Years 1, 2, and 3? Do not round intermediate calculations. Round your answers to the nearest dollar. Year 1 $ Year 2 $ Year 3 $ c. What is the additional Year 3 cash flow (i.e, the after-tax salvage and the return of working capital)? Do not round intermediate calculations. Round your answer to the nearest dollar. $ d. If the project's cost of capital is 14 %, what is the NPV of the project? Do not round intermediate calculations. Round your answer to the nearest dollar. $ Should the machine be purchased?

16. Broussard Skateboard's sales are expected to increase by 25% from $7.2 million in 2016 to $9.00 million in 2017. Its assets totaled $5 million at the end of 2016. Broussard is already at full capacity, so its assets must grow at the same rate as projected sales. At the end of 2016, current liabilities were $1.4 million, consisting of $450,000 of accounts payable, $500,000 of notes payable, and $450,000 of accruals. The after-tax profit margin is forecasted to be 3%, and the forecasted payout ratio is 60%. Use the AFN equation to forecast Broussard's additional funds needed for the coming year. Round your answer to the nearest dollar. Do not round intermediate calculations. $

17. Broussard Skateboard's sales are expected to increase by 25% from $8.4 million in 2016 to $10.50 million in 2017. Its assets totaled $6 million at the end of 2016. Broussard is already at full capacity, so its assets must grow at the same rate as projected sales. At the end of 2016, current liabilities were $1.4 million, consisting of $450,000 of accounts payable, $500,000 of notes payable, and $450,000 of accruals. The after-tax profit margin is forecasted to be 7%, and the forecasted payout ratio is 70%. What would be the additional funds needed? Do not round intermediate calculations. Round your answer to the nearest dollar. $ Assume that an otherwise identical firm had $7 million in total assets at the end of 2016. Broussard's capital intensity ratio (A0*/S0) is than the otherwise identical firm; therefore, Broussard is capital intensive - it would require increase in total assets to support the increase in sales.
Answer #1

Answer 13 - The after-tax net salvage value is $1,980,000

Reason - The book value of the equipment is 20% of its original cost since 80% was already depreciated.

Book Value = 22,000,000 * 20% = 4,400,000

However the equipment can be sold at a higher price of 7,700,000 and thus will result to a gain.

Gain = 7,700,000 - 4,400,000 = 3,300,000

Since the sale resulted to a gain, the Company is liable to a tax of 40% on the gain. So after the tax is paid, only 60% of the gain will be retained for the Company's disposal.

3,300,000 * 60% = 1,980,000

As per HOMEWORKLIB RULES first full question has been answered with explaination. Please provide other questions saperate to get answers for each of them.

If the solution helped, please give it a thumbs up. Thank you.